A restaurant has a dining room that is 62 ft by 55 ft. What is the area of the dining room? What is the perimeter of the dining room?
Found 2 solutions by rfer, JBarnum:
Answer by rfer(16322)

A=62*55
A=3410 sq ft
------
P=2*62+2*55
P=124+110
P=234 ft

Answer by JBarnum(2146)

A=L*W
L=62
W=55
62*55=3410


P=2(L+W)
P=2(62+55)
P=2(117)
P=234